OpenStudy (anonymous):

What is the simplified form of x+9/4 + x+5/4

OpenStudy (anonymous):

I understand 0% of this lesson... Help?

OpenStudy (anonymous):

Sorry, typo. x+9/x+1 + x-2/5

OpenStudy (anonymous):

oh ok thats better

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

Here are the choices:

OpenStudy (mathmale):

Look at the two given fractions which are being added together. What is the LCD of this set of 2 fractions? LCD = lowest common denominator.

OpenStudy (anonymous):

1?

OpenStudy (mathmale):

what does LCD mean to you? Hint: It's not 1.

OpenStudy (anonymous):

Lowest common denominator.

OpenStudy (mathmale):

Yes, and what does lowest common denominator mean in this context?

OpenStudy (anonymous):

Denominator is the lower part of the fraction.

OpenStudy (mathmale):

yes, and what does lowest common denominator mean? How do you find the LCD if you are given 2 fractions with different denoms. and are asked to add the fractions together?

OpenStudy (anonymous):

I'm not sure off the top of my head what LCD mean in this context.

OpenStudy (mathmale):

Better look up Lowest Common Denominator. Or, given 2/3 + 7/8, find the lcd and then add these fractions together.

OpenStudy (anonymous):

LCD means the least common multiple of two numbers. So, in this question, 5?

OpenStudy (mathmale):

Sorry, no. How did you get 5? Try again. what is the LCD of the 2 fractions, \[\frac{ 2 }{ 3 },\frac{ 7 }{ 8 }?\]

OpenStudy (anonymous):

Sorry, I ment the question I asked. The LCD of that question is 24.

OpenStudy (anonymous):

For mine it should be 5, right?

OpenStudy (mathmale):

all right. add the 2 fractions together, please.

OpenStudy (anonymous):

9/24, or 3/8?

OpenStudy (anonymous):

3/8 being the simplified form?

OpenStudy (mathmale):

are you telling me that the sum of 2/3 and 7/8 is 3/8?

OpenStudy (anonymous):

Yes. Is that wrong?

OpenStudy (mathmale):

definitely, I'm sorry to say. How long has it been since you last studied "LCD"?

OpenStudy (anonymous):

Since 7th grade. I'm in 10th now.

OpenStudy (anonymous):

Thats the last time it was reviewed.

OpenStudy (mathmale):

Seems as tho' you may need a review. Of course I can help with LCD's, but would prefer not to start from scratch. Since you have internet access, could you look up LCD?

OpenStudy (mathmale):

To combine two fractions, you must either have the same denom. in each, or know how to obtain and use the LCD.

OpenStudy (anonymous):

Okay

OpenStudy (mathmale):

1/3 - (4/3) = ?

OpenStudy (anonymous):

1?

OpenStudy (mathmale):

No, try again.

OpenStudy (anonymous):

:/

OpenStudy (mathmale):

sign error

OpenStudy (anonymous):

-1?

OpenStudy (mathmale):

Better. Right. find the lcd of 1/x and 2/y

OpenStudy (anonymous):

3?

OpenStudy (mathmale):

No. If you're interested, here's the URL of a web site that discusses LCDs and how to find them: https://www.khanacademy.org/math/pre-algebra/fractions-pre-alg/equivalent-fractions-pre-alg/v/finding-common-denominators

OpenStudy (mathmale):

I'd like to help you, but much prefer you do some reading about LCDs and how to find and use them before we continue. Much material is available on the Internet.

OpenStudy (anonymous):

I think I'll just find a different source of help. It might be googling the question since is 5 hours past my teachers hours.

OpenStudy (mathmale):

any source of info that works for you would be a good starting point.

OpenStudy (anonymous):

I have five other tests that need done before 10 tonight. My time, thats only two hours.

OpenStudy (mathmale):

i find the Internet to be a wonderful source of info on the spot. I very much sympathize, but also wonder how five tests could pile u p into one night. How long ago were you asked to do these assignments?

OpenStudy (anonymous):

the one I'm doing now needed sent in by the 24th of November.

OpenStudy (anonymous):

In short, I'm behind.

OpenStudy (mathmale):

Surely sounds like it. How did that happen? Do you schedule your study time?

OpenStudy (anonymous):

No. Personal issues murdered my sleep schedule and I haven't been able to make it to the online meetings in the mornings.

OpenStudy (mathmale):

Let's make the most of the scant time we have. are y ou or are you not going to do some quick research on LCDs on the 'Net? Exactly what do you hope to get from me, other than answers?

OpenStudy (anonymous):

I was hoping to get quick help with small explanations.

OpenStudy (anonymous):

Seeing as i asked this over a half hour ago, I assume I'm on the wrong site for that kind of help.

OpenStudy (mathmale):

Let's not be sarcastic. I object to having to explain concepts from scratch when you could reasonably be expected to have studied and practiced the basics.

OpenStudy (anonymous):

Like I said, my schedule's been tight due to personal issues. I would study the stuff harder if I had the time for it.

OpenStudy (mathmale):

You must have the same denominator in 2 fractions BEFORE you can add or subtract them. That's the principle. The next quesion is HOW DO YOU FIND THE LCD? what do you remember? I'm sorry about your having personal issues to deal with.

OpenStudy (mathmale):

Add:\[\frac{ 3 }{ 7 }+\frac{ 2 }{ 5 }\]

OpenStudy (mathmale):

to find the lcd, multiply the 7 and 5 together.

OpenStudy (anonymous):

3/35 + 2/35?

OpenStudy (anonymous):

21/35 + 14/35?

OpenStudy (mathmale):

The first fraction is 3/7. If you mult. the denom. by 7, you must also mult the numerator by 7. Do that now, please.

OpenStudy (mathmale):

Much b etter.

OpenStudy (anonymous):

Then simplify?

OpenStudy (mathmale):

Looking at the second fraction, whose den. is 5, mult both numerator and den. by 7 .

OpenStudy (anonymous):

14/35?

OpenStudy (mathmale):

\[\frac{ 3 }{ 7 }+\frac{ 2 }{ 5 }\]

OpenStudy (mathmale):

Yes. You've mult both num. and den. of the 2nd fraction by 7. I made a mistake on the first fract. Mult both its num. and den. by 5, please.

OpenStudy (anonymous):

10/25?

OpenStudy (mathmale):

5 times 3 is 10?

OpenStudy (anonymous):

Sorry, 15

OpenStudy (anonymous):

15/35

OpenStudy (mathmale):

better. Now you have \[\frac{ 15 }{ 35 }+\frac{ 14 }{ 35 }\]

OpenStudy (mathmale):

What is the LCD?

OpenStudy (anonymous):

the LCD is 35.

OpenStudy (mathmale):

Good. Now becasue you have the same den. in both fract., you can add the numerators to get the final answer (the sum)

OpenStudy (mathmale):

Do so, please.

OpenStudy (anonymous):

The simplified answer should be 29/35, right?

OpenStudy (mathmale):

Can't be simplified. Yes, the final answer is 29/35. Congrats.

OpenStudy (mathmale):

Now, find the LCD, use it and then add together \[\frac{ 3 }{ 7 }+\frac{ 2 }{ x }\]

OpenStudy (mathmale):

the LCD is ... ?

OpenStudy (anonymous):

7?

OpenStudy (mathmale):

How did you get that? Have you used multiplication here?

OpenStudy (anonymous):

7 * x = 7 since I don't know what x represents.

OpenStudy (mathmale):

In this case you MUST mult two quantities together. Yes, the LCD is 7x.

OpenStudy (anonymous):

Okay.

OpenStudy (mathmale):

Given \[\frac{ 3 }{ 7 }+\frac{ 2 }{ x }\]

OpenStudy (mathmale):

by what quantity must you mult. the num. and the den. of the first fract. ? why?

OpenStudy (anonymous):

Is this cross multiplication?

OpenStudy (mathmale):

No. focus on the 1st fract. Determine by what quant. you must mult. its num. and den.

OpenStudy (anonymous):

By x, right?

OpenStudy (mathmale):

Exactly. Do so, please.

OpenStudy (anonymous):

3x / 7x?

OpenStudy (mathmale):

very nice. By what quant. must you mult num. and den. of the 2nd fract?

OpenStudy (anonymous):

By x as well?

OpenStudy (mathmale):

No. What's the LCD here?\[\frac{ 3 }{ 7 }+\frac{ 2 }{ x }\]

OpenStudy (anonymous):

7x

OpenStudy (anonymous):

So, 2x/7x?

OpenStudy (mathmale):

Thus, mult the num and den of the 2nd fract by what quantity?

OpenStudy (anonymous):

By x.

OpenStudy (mathmale):

No, we've already used x to modify the 1st fract. Now we are modifying the 2nd fract.